OpenStudy (anonymous):

What is 5(3m - 4)^2 - 2(3m - 4) - 16 = 0, if u = 3m - 4

OpenStudy (anonymous):

I only got up to 5u^2 - 2u - 16 = 0

OpenStudy (anonymous):

i hate factoring, don't you?

OpenStudy (tiffanymak1996):

same as last time list out the factors

OpenStudy (tiffanymak1996):

of both the first and last terms.

OpenStudy (anonymous):

I don't really get it this time

OpenStudy (anonymous):

too many choices \[(5u+a)(u+b)\] i would solve the quadratic equation, then i would know what the factors are

OpenStudy (tiffanymak1996):

it's alright, the above would work too.

OpenStudy (tiffanymak1996):

5 only has 5x1

OpenStudy (anonymous):

Yes

OpenStudy (tiffanymak1996):

what about 16?

OpenStudy (anonymous):

yea, but 16 has a bunch 1, 16, 2, 8 4, 4

OpenStudy (anonymous):

and also 16, 1 8, 2 4, 4

OpenStudy (tiffanymak1996):

you try to find one pair that when times 5 the numbers are close

OpenStudy (anonymous):

and it is negative, so there are really twice as many ick ick ick

OpenStudy (anonymous):

I really don't know. I hate this stuff!

OpenStudy (tiffanymak1996):

especially when it's negative let's say using your brackets: (5u+a)(u+b)

OpenStudy (tiffanymak1996):

either 5b>a or a>5b both without the negative sign

OpenStudy (anonymous):

I don't get it

OpenStudy (tiffanymak1996):

i'll give an example

OpenStudy (tiffanymak1996):

say: x-3x-4

OpenStudy (tiffanymak1996):

-4=-4x1

OpenStudy (tiffanymak1996):

and -4+1 =-3

OpenStudy (tiffanymak1996):

do you get this?

OpenStudy (anonymous):

Yes

OpenStudy (tiffanymak1996):

okay, now let's work with -16 -16=-16x1 = -8x2 =-4x4 etc

OpenStudy (tiffanymak1996):

you multiply 5 into one number of the pairs

OpenStudy (anonymous):

ok

OpenStudy (tiffanymak1996):

try them, add them up with the negative sign, and try to find -2

OpenStudy (anonymous):

Ok. Thanks!

OpenStudy (tiffanymak1996):

you got it? gr8!

OpenStudy (tiffanymak1996):

so what's the answer?